OpenStudy (anonymous):

Pleaseeeee help Three quarts of a 24% iodine solution were mixed with a 52% solution to make a 40% iodine solution. How many quarts of the 52% solution were needed?

OpenStudy (kropot72):

Let q = number of quarts of final 40% iodine solution. Amount of iodine in 24% solution = 3 * 0.24 Amount of iodine in final 40% solution = q * 0.4 Amount of iodine in 52% solution = (q - 3) * 0.52 (3 * 0.24) + 0.52(q - 3) = 0.4q Now you just need to solve the last equation to find q (the number of quarts of final 40% iodine solution) and subtract 3 from the result to find the number of quarts of the 52% solution that were needed.